1.A metal object at 38.0 oC is heated by gaining 11.5 kJ of heat
from the environment.
The heat capacity, C, of the metal object is 1.55 kJ/ oC.
(Note: NOT the specific heat, which has units of J/ g oC)
  
What is the final T?

2.A metal cube loses 3.358 kJ of heat. (i.e., q = -3.358 kJ)
The metal has mass of 135.6 grams. The initial temperature of the metal before the heat loss is is 158.0 oC, the final temperature (after heat loss) is 35.0 oC. What is the specific heat, c, of the metal in units of J/g oC?

3.A metal object at 54.0 oC is heated by gaining 30.2 kJ of heat
from the environment.
The heat capacity, C, of the metal object is 6.30 kJ/ oC.
(Note: NOT the specific heat, which has units of J/ g oC)
  
What is the final T?

4.An object at 50.0 oC is heated to 87.0 oC by gaining 7.75 kJ of heat
from the environment.
What is the heat capacity of the entire object in units of kJ/ oC?  
(NOT the specific heat, which has units of J/ g oC)
Answer #1

Lo Initial tent. ( T ) = 38.0°c a = 115 KJ = 11500T Tmetal= 1.55 KIloc = 155oT T2=P a = Thetal & OT 115005= 1550 Jlext T2-38TIF 50.0°C Ta = 87.00 9 = 7.75 KJ 29500J Cobject = ? a = Tobject & OT 77500 J- Cobject X (07.0-50.0) Cobject = 77580 - 20g4.5